Shift Summary for KTeV - 9/4/99 DAY (Sat.)

Earl C. Swallow (earls@elmhurst.edu)
Sat, 4 Sep 1999 15:37:22 -0500 (CDT)

Shift Summary for DAY shift, 09/04/99 (Sat.)

On Shift: Elizabeth Turner & Earl Swallow

Hours of beam available at all intensities: 8.1 (out of 8.5)
Hours of beam available at nominal intensity: 8.1
Hours of E832 data (special runs not included): 7.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

2 (count'em 2!) tape change & several KFRENDS +
accelerator energy conservation & aborts.


========================================================================
Runs ended this shift:
14426 - BEAM_832
14427 - MEAN_832

Tapes brought to FCC this shift:
QK(C,S,K)378-380, QKB133 (Runs 14424, 14425)
QK(C,S,K)381-383, QKB134 (Run 14426)
QK(C,S,K)384-386, QKB135 (Run 14427)

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

VERY Quiet shift with steady data taking -- HIP HIP HOORAY!
(2 tape changes in one shift!)

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14426 07:00 Shift Begins.
Tapes are almost full.
There is an intermittant VolMon2 error for Rack 147,
NIM crate 1 (RR174-NIM1). The +6 Volt supply
sometimes reads +3.8 - +4.6 Volts.
14426 07:11 END Run 14426 - BEAM_832.
Remove tapes: QK(C,S,K)381-383, QKB134.
14427 1 07:19 BEGIN Run 14427 - BEAM_832.
Mounted Tapes: QK(C,S,K)384-386, QKB135. NOTE: Both
members of Shift Crew checked tape numbers before
eating lunch.
07:40 Mounted ONLINE SPLIT tape: QKPE14 --> KTEV3/tps2d4.

08:00 JasonL has come in to investigate the RR174-NIM1
"problem." His conclusion: that particular crate is
not currently in regular use -- some modules in it
are for TRD trigger, and he uses one module for an
occasional calibration procedure. He posts a
"don't sweat it" note on SLOW-DAQ terminal.

163 11:00 Scaler ratios out of bounds. Looks like they
weren't read out. Fixes itself.

12:00 Intensity down to ~5.7E12 -- some problems
balancing beam for NUMI target test and TeVatron
fixed target. Max in MCR is attending to it.

252 13:05 Max from MCR calls to say they found and "bypassed"
a power supply problem between Linac and Booster.
This caused intensity to jump up to ~7.5E12, but
they are bringing it back down to match our
request.

14427 341 15:05 END Run 14427 - BEAM_832.

14428 1 15:10 BEGIN Run 14428 - BEAM_832.

15:30 Shift ends with continued data taking.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

--
===============================================================================
Earl C. Swallow                           email:  earls@elmhurst.edu
Theophil W. Mueller Professor              (or)   earls@fnal.gov
   and Chairperson
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu/~earls
===============================================================================
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)